Solve 2= 3 square root of x - 5

1
Expert's answer
2021-03-18T06:40:41-0400

2 = "\sqrt[3]{x-5}"

23 = ("\sqrt[3]{x-5}" )3

8 = x - 5

8 + 5 = x

x = 13

The answer is x = 13.